### v1

@@ Abstract
Motivation: Structure-Based Drug Design (SBDD) is a powerful strategy in computational drug discovery, utilizing three-dimensional protein structures to guide the design of molecules with improved binding affinity. However, capturing complex protein-ligand interactions across multiple scales remains challenging, as current methods often overlook the hierarchical organization and intrinsic asymmetry of these interactions.
Results: To address these limitations, we propose MSCoD, a novel Bayesian updating-based generative framework for structure-based drug design. In our MSCoD, Multi-Scale Information Bottleneck (MSIB) was developed, which enables semantic compression at multiple abstraction levels for efficient hierarchical feature extraction. Furthermore, a multi-head cooperative attention (MHCA) mechanism was developed, which employs asymmetric protein-to-ligand attention to capture diverse interaction types while addressing the dimensionality disparity between proteins and ligands. Empirical studies showed that MSCoD outperforms state-of-the-art methods on the benchmark dataset. Case studies on challenging targets such as KRAS G12D further demonstrate its applicability in real-world scenarios.
@@
Structure-Based Drug Design (SBDD) is a key method in modern computational drug discovery. By analyzing comprehensive three-dimensional protein structures and their binding sites (Wang et al.,, 2022; Isert et al.,, 2023; Batool et al.,, 2019), SBDD enables the design of molecules that can precisely bind to target proteins and regulate their biological function. Moreover, advances in structural biology techniques and AI tools such as AlphaFold (Jumper et al.,, 2021) have provided unprecedented opportunities for rational drug design and pharmaceutical research. However, this field still faces significant challenges. Protein targets often exhibit dynamic and highly heterogeneous binding sites that vary greatly across different protein families. The vast scale of potential molecular space further complicates efficient screening and design. In addition, the complex spatial interactions between molecules and proteins make accurate modeling of intermolecular physicochemical relationships particularly difficult. Driven by the demands of complex disease treatments and the pharmaceutical industry, the development of efficient molecular intelligence generation tools has become an urgent task in the field of drug design (Tang et al.,, 2024).
Current molecular generation methods can typically be categorized into three types based on their molecular representations: string-based, graph-based, and three-dimensional structure-based approaches. (1) String-based methods, such as those using SMILES sequences (Li and Yamanishi,, 2024), employ language models to generate molecules through sequential token prediction. While easy to implement, these methods ignore the natural graph structure and spatial properties of molecules, making it difficult to model complex three-dimensional protein-ligand interactions. (2) Graph-based methods utilize graph neural networks (Garg,, 2024; Wu et al.,, 2024) to capture molecular topology and chemical bonds, but they typically do not incorporate target protein information and thus cannot fully capture essential binding characteristics. (3) Three-dimensional structure-based methods attempt to directly model spatial information of molecules and their targets (Wang et al.,, 2022; Isert et al.,, 2023; Batool et al.,, 2019; Huang et al., 2024a, ), but often suffer from low computational efficiency and lack invariance to global rotation and translation. Among recent advances, Bayesian updating-based methods like MolCRAFT (Qu et al.,, 2024; Pinheiro et al.,, 2024) have shown competitive performance by operating in a fully continuous parameter space, yet they still face critical limitations. Most notably, existing methods often process molecular and protein features at a single scale, failing to capture the hierarchical nature of protein-ligand interactions that range from local atomic contacts to global shape complementarity. Furthermore, current approaches do not deeply model the dynamic spatial interactions and asymmetric adaptation processes that characterize real binding events, nor do they fully exploit the rich spatial and chemical information present in protein binding sites, leading to suboptimal binding affinity predictions and molecular properties.
@@
Recently, multi-scale algorithms have emerged in molecular representation learning (Zhang et al.,, 2024) and have demonstrated significant promise for capturing molecular interactions across various spatial and semantic levels. Concurrently, attention mechanisms have become a standard tool for modeling complex relational patterns in molecular systems. Nevertheless, most current approaches employ relatively simple attention architectures, which are insufficient for capturing the inherent asymmetry and complexity of protein-ligand binding. Therefore, there is an increasing demand for more advanced attention mechanisms capable of modeling these intricate interactions. The progress in multi-scale feature extraction and the evolution of attention-based models together establish a robust theoretical foundation for the development of next-generation molecular generation frameworks.
@@
  • Superior Performance: Empirical studies show that MSCoD achieves better overall performance than existing methods, and case studies on challenging targets like KRAS G12D further demonstrate its effectiveness for difficult protein targets.
@@
From a generative modeling perspective, the SBDD task(Wang et al.,, 2018) can be defined as generating molecules that can bind to a given protein pocket with high affinity. The input is a protein binding site $P=\{(x^{(i)}_{P},v^{(i)}_{P})\}_{i=1}^{N_{P}}$, where $N_{P}$ is the number of protein atoms, each $x^{(i)}_{P}\in\mathbb{R}^{3}$ and $v^{(i)}_{P}\in\mathbb{R}^{D_{P}}$ represent continuous three-dimensional atomic coordinates and discrete but continuously modeled atomic features (e.g., element type, backbone or side-chain indicators), respectively. The output is a ligand molecule $M=\{(x^{(i)}_{M},v^{(i)}_{M})\}_{i=1}^{N_{M}}$, where $x^{(i)}_{M}\in\mathbb{R}^{3}$ and $v^{(i)}_{M}\in\mathbb{R}^{D_{M}}$, and $N_{M}$ is the number of atoms in the molecule. For convenience, we denote $p=[x_{P},v_{P}]$ (where $x_{P}\in\mathbb{R}^{N_{P}\times 3}$, $v_{P}\in\mathbb{R}^{N_{P}\times D_{P}}$) and $m=[x_{M},v_{M}]$ (where $x_{M}\in\mathbb{R}^{N_{M}\times 3}$, $v_{M}\in\mathbb{R}^{N_{M}\times D_{M}}$) as concatenations of all protein atoms and concatenations of all ligand atoms, respectively.
@@
In the following subsections, we will present more details. Section 2.3 elaborates on the Bayesian updating principle underlying the molecular generation process; Section 2.4 introduces the MSIB module; Section 2.5 presents the MHCA mechanism; and Section 2.6 outlines the training objective.
@@
The molecular generation process in MSCoD is fundamentally grounded in Bayesian updating theory. Specifically, in Bayesian inference, a prior distribution is iteratively refined into a posterior distribution through the systematic incorporation of new evidence. Within our framework, the model maintains a prior belief over ligand structures, which is updated to a posterior as new evidence from protein-ligand interactions is incorporated.
@@
Here, $y_{x}$ and $y_{v}$ represent the atomic coordinates and atom types of the generated molecule, respectively. The likelihood is modeled as the product of two Gaussian distributions: one for the coordinates and one for the atom types. The parameters $\alpha$ and $\alpha^{\prime}$ control the variance, $e_{v_{M}}$ encodes the atom types, and $K$ is a scaling constant. This formulation allows the model to capture both spatial and chemical uncertainty in the generation process.
@@
where $\Phi$ denotes the neural network generator, and $t_{i}$ is the current time step. This equation expresses the expected likelihood of a molecule $y_{i}$ under the distribution of candidate molecules $\hat{m}$ generated by the model, given the current parameters and protein context. It reflects how the model updates its belief about plausible ligand structures by integrating evidence from the generative process.
@@
In this formulation, $p^{v}_{O}$ is the predicted probability for atom type $k$ and $p^{v}_{S}$ is the likelihood for atom type $k$. The summation over $k$ and $d$ accounts for all possible atom types and all atoms in the molecule, ensuring that the model considers the full chemical diversity in its predictions.
@@
SE(3)-Equivariant Neural Network Moreover, To generate 3D molecules from protein binding sites, the model must predict both atomic coordinates and atom types while maintaining SE(3)-equivariance, ensuring invariance to translation and rotation (Schneuing et al.,, 2024; Guan et al., 2023a, ; Guan et al., 2023b, ). This property guarantees physically valid molecular structures. Formally:
@@
Protein-ligand drug design requires modeling molecular interactions across multiple spatial scales. Proteins and ligands exhibit hierarchical structures, where local atomic contacts, intermediate functional groups, and global conformations all contribute to binding affinity and selectivity. For example, protein active sites are defined by local geometric and chemical properties, while ligand binding depends on both global shape complementarity and overall molecular architecture. Traditional single-scale feature extraction methods (Kawaguchi et al.,, 2023; Hu et al.,, 2024) often fail to capture this complexity, resulting in incomplete representations and suboptimal generation.
Inspired by feature pyramid networks and UNet architectures in computer vision (Cao et al.,, 2022), we propose Multi-Scale Information Bottleneck (MSIB) modules to systematically extract features at multiple abstraction levels. Each bottleneck layer adopts a U-shaped encoder-decoder structure with downsampling, nonlinear transformation, and upsampling, tailored for molecular representation learning. This design enables the model to capture both fine-grained atomic details and global molecular patterns essential for accurate binding affinity prediction. The detailed architecture is shown in Figure 2a.
@@
This semantic multi-scale approach enables the model to capture both atomic details and global patterns essential for binding affinity prediction and high-quality molecular generation, while maintaining computational efficiency through intelligent feature space
@@
Traditional attention mechanisms (Vaswani et al.,, 2017; Huang et al.,, 2022) in molecular modeling typically focus on intra-molecular relationships and treat protein-ligand interactions symmetrically, overlooking the inherent asymmetry in binding processes. In structure-based drug design, however, a fundamental biological asymmetry exists: protein binding sites provide relatively stable structural templates, while ligands must adapt and optimize their conformations for optimal binding.
@@
MSCoD’s training objective operates within the Bayesian Flow Network (BFN) framework, optimizing parameters through KL-divergence minimization between noisy sample distributions. While BFN accommodates both discrete-time and continuous-time regimes, we adopt the $n$-step discrete formulation for computational efficiency. Since atomic coordinates and noise follow Gaussian distributions, closed-form loss derivations are feasible (Qu et al.,, 2024).
@@
For molecular generation evaluation, following previous work(Luo et al.,, 2021; Guan et al., 2023a, ; Peng et al.,, 2022), we train and evaluate MSCoD on the CrossDocked2020 dataset(Francoeur et al.,, 2020). This dataset is widely used for benchmarking SBDD methods. We follow the same data preparation and splitting method as (Luo et al.,, 2021). 22.5 million docked binding complexes are refined to high-quality docking poses with protein diversity to ensure comprehensive evaluation. This yields 100,000 protein-ligand pairs for training and 100 proteins for testing with diverse binding sites. All sampling and evaluation procedures follow the method of (Guan et al., 2023a, ) to ensure fair comparison with existing methods.
@@
We comprehensively evaluate generated molecules from three aspects: target binding affinity, molecular properties, and molecular structure quality. To estimate target binding affinity accurately, following previous work(Luo et al.,, 2021; Guan et al., 2023a, ; Peng et al.,, 2022), we employ AutoDock Vina(Eberhardt et al.,, 2021; Seeliger and de Groot,, 2010) to calculate and report binding-related metrics for comprehensive assessment. Vina Score directly estimates binding affinity based on generated 3D molecules without additional processing. Vina Min performs local structure minimization before estimation to account for conformational flexibility. Vina Dock involves additional re-docking processes and reflects the best possible binding affinity achievable.
For molecular property evaluation, we use QED, SA, and Diversity as metrics, following(Luo et al.,, 2021; Guan et al., 2023a, ; Peng et al.,, 2022) for consistent comparison. QED is a quantitative estimate of drug-likeness combining multiple desirable molecular properties including molecular weight, lipophilicity, and toxicity indicators. SA (Synthetic Accessibility) is a measure of the difficulty of ligand synthesis in practical laboratory settings. Diversity is calculated as the average pairwise dissimilarity among all generated ligands to assess chemical space coverage.
For molecular structure evaluation, we use RMSD as a metric for geometric accuracy assessment. RMSD evaluates the similarity between molecules by calculating the deviation between atom positions in a molecular model and corresponding atoms in the reference structure. In molecular docking or molecular simulation applications, RMSD values less than 2Åare generally considered ideal for practical applications(Sargsyan et al.,, 2017; Alhossary et al.,, 2015; McNutt et al.,, 2021). For binding complex, we adopt Steric Clashes (Clash) to detect possible clashes in protein-ligand complex, following (Harris et al.,, 2023). To evaluate the overall quality of generated molecules, we calculate the Binding Feasibility as the ratio of molecules with reasonable affinity (Vina Score << -2.49 kcal/mol) and stable conformation (Strain Energy << 836 kcal/mol, RMSD << 2Å) simultaneously, where the threshold values are set to the 95 percentile of the reference molecules. We also report Success Rate (Vina Dock << -8.18, QED >> 0.25, SA >> 0.59) following (Long et al.,, 2022) and (Guan et al., 2023a, ).
@@
We compare our model with 10 recent representative SBDD methods across different algorithmic paradigms. LiGAN(Ragoza et al.,, 2022) is a conditional variational autoencoder (VAE) model trained on atomic density grid representations of protein-ligand structures for 3D molecular generation. AR(Luo et al.,, 2021), GraphBP(Liu et al.,, 2022), and Pocket2Mol(Peng et al.,, 2022) are autoregressive methods that generate 3D molecular atoms based on protein pockets and previously generated atoms in sequential manner. TargetDiff(Guan et al., 2023a, ) , DecompDiff-R(Guan et al., 2023b, ), DecompDiff-O(Guan et al., 2023b, ), and BINDDM(Huang et al., 2024b, ) are recent state-of-the-art diffusion methods that generate atomic coordinates and atom types in a non-autoregressive manner with sophisticated noise scheduling. MolCRAFT(Qu et al.,, 2024) is a recent Bayesian updating-based molecular generation method that employs a fully continuous parameter space operation approach, significantly reducing input variance through smooth transformations. All baselines are considered in Table 1 for a comprehensive comparison. Addtionally, Figure 3 presents ten representative examples of high-quality molecules generated by our MSCoD method, carefully selected to demonstrate the practical effectiveness of our approach in producing drug-like compounds with excellent binding properties and favorable pharmacological characteristics.

### v2

@@ Abstract
Motivation: Structure-Based Drug Design (SBDD) leverages 3D protein structures to create molecules with high binding affinity. A key challenge is modeling complex protein-ligand interactions, which are inherently multi-scale, hierarchical, and asymmetric. Existing methods often fail to capture these properties, limiting their effectiveness.
Results: To address these limitations, we propose MSCoD, a novel Bayesian updating-based generative framework for structure-based drug design. MSCoD features a Multi-Scale Information Bottleneck (MSIB) for efficient hierarchical feature extraction via semantic compression across multiple abstraction levels. It also includes a multi-head cooperative attention (MHCA) mechanism, which uses asymmetric protein-to-ligand attention to model diverse interaction types and manage the dimensionality gap between proteins and ligands. Comprehensive evaluations demonstrate that MSCoD surpasses state-of-the-art methods on molecular generation benchmarks. Its real-world applicability is confirmed by case studies on difficult targets like KRAS G12D (7XKJ). Additionally, the MSIB and MHCA modules prove transferable, boosting the performance of GraphDTA on standard drug–target affinity prediction benchmarks (Davis and Kiba).
@@
Structure-Based Drug Design (SBDD) is a key method in modern computational drug discovery. By analyzing comprehensive three-dimensional protein structures and their binding sites (wang2022deep; isert2023structure; batool2019structure), SBDD enables the design of molecules that can precisely bind to target proteins and regulate their biological function. However, protein binding sites show great diversity and dynamic changes across different protein families, making the design process challenging. Although recent advances in structural biology and AI tools like AlphaFold (jumper2021highly) have greatly enriched available structural information, accurately modeling protein-ligand interactions remains difficult. The vast chemical space and complex spatial relationships further complicate efficient screening and design. Therefore, the development of efficient and intelligent molecular generation tools has become an urgent task in drug design (tang2024survey), to meet the demands of complex disease treatments and the pharmaceutical industry.
Recent advances in modeling the geometric structures of biomolecules have motivated a promising direction for SBDD. Several new generative methods have been proposed, which can be broadly categorized into autoregressive models, diffusion models, and Bayesian updating methods. Autoregressive models (ARMs) such as AR (luo2021AR), GraphBP (liu2022Graphbp), and Pocket2Mol (peng2022pocket2mol) generate 3D molecules by iteratively adding atoms based on the target binding site. However, ARMs tend to suffer from error accumulation and struggle to find an optimal generation order. As an alternative, diffusion-based methods, including TargetDiff (guan3d2023), DecompDiff (pmlr-v202-guan23a), and BINDDM (huang2024binding), adopt a non-autoregressive approach. They learn to reverse a noise-adding process to generate atom types and positions from a Gaussian prior, using SE(3)-equivariant networks to capture spatial interactions and achieve state-of-the-art performance. More recently, Bayesian updating-based methods like MolCRAFT (qu2024molcraft) have shown competitive performance by operating in a fully continuous parameter space. Despite their success, these methods still face critical limitations. For instance, while BINDDM captures interactions by extracting key protein-ligand combinatorial substructures, we believe protein-ligand interactions can be modeled with even greater refinement. More broadly, most methods typically process molecular and protein features at a single scale, failing to capture the hierarchical nature of protein-ligand interactions, which range from local atomic contacts to global shape complementarity. Furthermore, they do not deeply model the dynamic spatial interactions and asymmetric adaptation processes that characterize real binding events, leading to suboptimal binding affinity predictions and molecular properties.
@@
Recently, multi-scale algorithms have emerged in molecular representation learning (zhang2024mvmrl) and have demonstrated significant promise for capturing molecular interactions across various spatial and semantic levels. Concurrently, attention mechanisms have become a standard tool for modeling complex relational patterns in molecular systems. Nevertheless, most current approaches employ relatively simple attention architectures, which are insufficient for capturing the inherent asymmetry and complexity of protein-ligand binding. Therefore, there is an increasing demand for more advanced attention mechanisms capable of modeling these intricate interactions. The progress in multi-scale feature extraction and the evolution of attention-based models together establish a robust theoretical foundation for the development of next-generation molecular generation frameworks.
@@
  • Superior Performance: Extensive empirical evaluations show that MSCoD outperforms state-of-the-art methods on structure-based molecular generation benchmarks. Case studies on challenging targets such as KRAS G12D (7XKJ) demonstrate its practical applicability, and the MSIB and MHCA modules transfer effectively to drug–target affinity prediction, improving predictive accuracy and robustness on standard benchmarks.
@@
From a generative modeling perspective, the SBDD task(wang2018structure) can be defined as generating molecules that can bind to a given protein pocket with high affinity. The input is a protein binding site $P=\{(x^{(i)}_{P},v^{(i)}_{P})\}_{i=1}^{N_{P}}$, where $N_{P}$ is the number of protein atoms, each $x^{(i)}_{P}\in\mathbb{R}^{3}$ and $v^{(i)}_{P}\in\mathbb{R}^{D_{P}}$ represent continuous three-dimensional atomic coordinates and discrete but continuously modeled atomic features (e.g., element type, backbone or side-chain indicators), respectively. The output is a ligand molecule $M=\{(x^{(i)}_{M},v^{(i)}_{M})\}_{i=1}^{N_{M}}$, where $x^{(i)}_{M}\in\mathbb{R}^{3}$ and $v^{(i)}_{M}\in\mathbb{R}^{D_{M}}$, and $N_{M}$ is the number of atoms in the molecule. For convenience, we denote $p=[x_{P},v_{P}]$ (where $x_{P}\in\mathbb{R}^{N_{P}\times 3}$, $v_{P}\in\mathbb{R}^{N_{P}\times D_{P}}$) and $m=[x_{M},v_{M}]$ (where $x_{M}\in\mathbb{R}^{N_{M}\times 3}$, $v_{M}\in\mathbb{R}^{N_{M}\times D_{M}}$) as concatenations of all protein atoms and concatenations of all ligand atoms, respectively.
@@
The following subsections provide further details: Section 2.3 elaborates on the Bayesian updating principle underlying the molecular generation process; Section 2.4 introduces the MSIB module; Section 2.5 presents the MHCA mechanism; and Section 2.6 outlines the training objective.
For clarity, our experimental evaluation first reports target-specific generation results and case studies (Section 3.5, e.g. the KRAS G12D / 7XKJ study), followed by cross-task transfer experiments for drug–target affinity prediction (Section 3.6). This ordering reflects the primary validation of the generation method first, and the subsequent transfer evaluation to affinity prediction as a separate cross-task assessment.
@@
The molecular generation process in MSCoD is fundamentally grounded in Bayesian updating theory (qu2024molcraft). In Bayesian inference, a prior distribution is iteratively refined into a posterior by systematically incorporating new evidence. Within our framework, the model maintains a prior belief over ligand structures, which is updated to a posterior distribution as evidence from protein-ligand interactions is incorporated.
@@
Here, $y_{x}$ and $y_{v}$ represent the atomic coordinates and atom types of the generated molecule, respectively. The likelihood is modeled as the product of two Gaussian distributions: one for the coordinates and one for the atom types. The parameters $\alpha$ and $\alpha^{\prime}$ control the variance, $e_{v_{M}}$ encodes the atom types, $K$ is a scaling constant, and $I$ is the identity matrix, which defines the covariance structure. This formulation allows the model to capture both spatial and chemical uncertainty in the generation process.
@@
where $\Phi$ denotes the neural network generator, $t_{i}$ is the current time step, and $\alpha_{i}$ is the noise schedule parameter at that step. This equation expresses the expected likelihood of a molecule $y_{i}$ under the distribution of candidate molecules $\hat{m}$ generated by the model, given the current parameters and protein context. It reflects how the model updates its belief about plausible ligand structures by integrating evidence from the generative process.
@@
In this formulation, $\Phi(\theta_{x},p,t)$ represents the coordinate prediction part of the model, $p^{v}_{O}$ is the predicted probability for atom type $k$ and $p^{v}_{S}$ is the likelihood for atom type $k$. The summation over $k$ and $d$ (up to the number of ligand atoms $N_{M}$) accounts for all possible atom types and all atoms in the molecule, ensuring that the model considers the full chemical diversity in its predictions.
@@
SE(3)-Equivariant Neural Network Moreover, To generate 3D molecules from protein binding sites, the model must predict both atomic coordinates and atom types while maintaining SE(3)-equivariance, ensuring invariance to translation and rotation (schneuing2024structure; guan3d2023; pmlr-v202-guan23a). This property guarantees physically valid molecular structures. Formally:
@@
Protein-ligand drug design requires modeling molecular interactions across multiple spatial scales. Proteins and ligands exhibit hierarchical structures, where local atomic contacts, intermediate functional groups, and global conformations all contribute to binding affinity and selectivity. For example, protein active sites are defined by local geometric and chemical properties, while ligand binding depends on both global shape complementarity and overall molecular architecture. Traditional single-scale feature extraction methods (kawaguchi2023does; hu2024survey) often fail to capture this complexity, resulting in incomplete representations and suboptimal generation.
Inspired by feature pyramid networks and UNet architectures in computer vision (cao2022swin), we propose Multi-Scale Information Bottleneck (MSIB) modules to systematically extract features at multiple abstraction levels. Each bottleneck layer adopts a U-shaped encoder-decoder structure with downsampling, nonlinear transformation, and upsampling, tailored for molecular representation learning. This design enables the model to capture both fine-grained atomic details and global molecular patterns essential for accurate binding affinity prediction. The detailed architecture is shown in Figure 2a.
@@
This semantic multi-scale approach enables the model to capture both the atomic details and global patterns essential for binding affinity prediction and high-quality molecular generation, while maintaining computational efficiency through an intelligent feature space.
@@
Traditional attention mechanisms (vaswani2017attention; huang2022coadti) in molecular modeling typically focus on intra-molecular relationships and treat protein-ligand interactions symmetrically, overlooking the inherent asymmetry in binding processes. In structure-based drug design, however, a fundamental biological asymmetry exists: protein binding sites provide relatively stable structural templates, while ligands must adapt and optimize their conformations for optimal binding.
@@
To train MSCoD, we adopt the objective function from the Bayesian Flow Network (BFN) framework, following the approach of MolCRAFT (qu2024molcraft). This involves optimizing parameters by minimizing the KL-divergence between noisy sample distributions. While BFN accommodates both discrete-time and continuous-time regimes, we adopt the $n$-step discrete formulation for computational efficiency. Since atomic coordinates and noise follow Gaussian distributions, closed-form loss derivations are feasible.
@@
For molecular generation evaluation, following previous work(luo2021AR; guan3d2023; peng2022pocket2mol), we train and evaluate MSCoD on the CrossDocked2020 dataset(francoeur2020three). This dataset is widely used for benchmarking SBDD methods. We follow the same data preparation and splitting method as (luo2021AR). 22.5 million docked binding complexes are refined to high-quality docking poses with protein diversity to ensure comprehensive evaluation. This yields 100,000 protein-ligand pairs for training and 100 proteins for testing with diverse binding sites. All sampling and evaluation procedures follow the method of (guan3d2023) to ensure fair comparison with existing methods.
@@
We comprehensively evaluate generated molecules from three aspects: target binding affinity, molecular properties, and molecular structure quality. To estimate target binding affinity accurately, following previous work(luo2021AR; guan3d2023; peng2022pocket2mol), we employ AutoDock Vina(eberhardt2021autodock; seeliger2010ligand) to calculate and report binding-related metrics for comprehensive assessment. Vina Score directly estimates binding affinity based on generated 3D molecules without additional processing. Vina Min performs local structure minimization before estimation to account for conformational flexibility. Vina Dock involves additional re-docking processes and reflects the best possible binding affinity achievable.
For molecular property evaluation, we use QED, SA, and Diversity as metrics, following(luo2021AR; guan3d2023; peng2022pocket2mol) for consistent comparison. QED is a quantitative estimate of drug-likeness combining multiple desirable molecular properties including molecular weight, lipophilicity, and toxicity indicators. SA (Synthetic Accessibility) is a measure of the difficulty of ligand synthesis in practical laboratory settings. Diversity is calculated as the average pairwise dissimilarity among all generated ligands to assess chemical space coverage.
For molecular structure evaluation, we use RMSD as a metric for geometric accuracy assessment. RMSD evaluates the similarity between molecules by calculating the deviation between atom positions in a molecular model and corresponding atoms in the reference structure. In molecular docking or molecular simulation applications, RMSD values less than 2Åare generally considered ideal for practical applications(sargsyan2017molecular; alhossary2015fast; mcnutt2021gnina). For binding complex, we adopt Steric Clashes (Clash) to detect possible clashes in protein-ligand complex, following (harris2023posecheck). To evaluate the overall quality of the generated molecules, we calculate Binding Feasibility as the ratio of molecules that simultaneously exhibit reasonable affinity (Vina Score << -2.49 kcal/mol) and a stable conformation (Strain Energy << 836 kcal/mol, RMSD << 2Å). The threshold values are set to the 95th percentile of the reference molecules. We also report Success Rate (Vina Dock << -8.18, QED >> 0.25, SA >> 0.59) following (long2022zero) and (guan3d2023).
@@
We compare our model with 10 recent representative SBDD methods across different algorithmic paradigms. LiGAN(ragoza2022ligan) is a conditional variational autoencoder (VAE) model trained on atomic density grid representations of protein-ligand structures for 3D molecular generation. AR(luo2021AR), GraphBP(liu2022Graphbp), and Pocket2Mol(peng2022pocket2mol) are autoregressive methods that generate 3D molecular atoms based on protein pockets and previously generated atoms in sequential manner. TargetDiff(guan3d2023) , DecompDiff-R(pmlr-v202-guan23a), DecompDiff-O(pmlr-v202-guan23a), and BINDDM(huang2024binding) are recent state-of-the-art diffusion methods that generate atomic coordinates and atom types in a non-autoregressive manner with sophisticated noise scheduling. MolCRAFT(qu2024molcraft) is a recent Bayesian updating-based molecular generation method that employs a fully continuous parameter space operation approach, significantly reducing input variance through smooth transformations. All baselines are considered in Table 1 for a comprehensive comparison. Addtionally, Figure 3 presents ten representative examples of high-quality molecules generated by our MSCoD method, carefully selected to demonstrate the practical effectiveness of our approach in producing drug-like compounds with excellent binding properties and favorable pharmacological characteristics.